The measure of an angle is 20 less than 3 times its supplement. Find the measure of the angle.
Answer by rapaljer(4671)

Let x = the angle
90 - x = the complement
180 - x= the supplement


x= 3(180 - x)-20
x = 540 - 3x-20

Add + 3x to each side:
4x = 520
x= 130 degrees = the angle
180 - x= 180 - 130 = 50 degrees = supplement.

Check:
The angle is 20 degees less than 3 times the supplement
130 = 150 - 20

It checks!!!

By the way, it's always good to check these problems, because the first time I worked this I missed it!! I found the error after I checked it.

Dr. Robert J. Rapalje, Retired